# Watermarking Recommender Systems

### v1

@@ Abstract.
Recommender systems embody significant commercial value and represent crucial intellectual property. However, the integrity of these systems is constantly challenged by malicious actors seeking to steal their underlying models. Safeguarding against such threats is paramount to upholding the rights and interests of the model owner. While model watermarking has emerged as a potent defense mechanism in various domains, its direct application to recommender systems remains unexplored and non-trivial. In this paper, we address this gap by introducing Autoregressive Out-of-distribution Watermarking (AOW), a novel technique tailored specifically for recommender systems. Our approach entails selecting an initial item and querying it through the oracle model, followed by the selection of subsequent items with small prediction scores. This iterative process generates a watermark sequence autoregressively, which is then ingrained into the model’s memory through training. To assess the efficacy of the watermark, the model is tasked with predicting the subsequent item given a truncated watermark sequence. Through extensive experimentation and analysis, we demonstrate the superior performance and robust properties of AOW. Notably, our watermarking technique exhibits high-confidence extraction capabilities and maintains effectiveness even in the face of distillation and fine-tuning processes. [^1] Accepted by CIKM 2024.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Data miningccs: Computing methodologies Neural networks
@@ 1. Introduction
Machine learning models have become a necessary component in the modern information society. One of the most representative machine learning product is the recommender system (rendle2012bpr, ; zhang2019deep, ; he2020lightgcn, ; wu2022graph, ; long2024physical, ). It has been widely deployed in various domains including health care (yue2021overview, ), finance (sharaf2022survey, ), e-commerce (schafer2001commerce, ), social platforms (fan2019graph, ), etc. It plays a crucial role in improving user experience and gaining profits for the service provider. As highly commercialized products, many of them involve certain intellectual property and copyright as well as business secrets. Some recommender systems are developed by the service provider themselves, where cutting-edge techniques might have been adopted to outperform competitors. In this case, model theft and model leakage are intolerable. Some other recommender systems are developed by a third party at the request of the service provider, where the illegal use and redistribution of the model are typically prohibited. Thus, it is necessary to take protective measures to ensure the recommender systems are not being used for these malicious aims.
@@
To the best of our knowledge, there are a limited number of works addressing the above issue on recommender systems. Zhang et al. (zhang2024defense, ) proposed an optimization-based method against model extraction attacks on recommender systems. They train the target model with an additional loss that can cause malfunctions to the attacker’s model. However, this method inevitably degrades the performance of the target model. As a commercial product, any degradation in performance for the recommender system may lead to a huge profit loss. Thus, it is necessary to design a protection method without sacrificing the model utility.
@@
One promising defense strategy in this scenario is the model watermarking. It has been extensively studied in computer vision (uchida2017embedding, ; chen2019deepmarks, ; wang2020watermarking, ; zhang2018protecting, ; adi2018turning, ; guo2018watermarking, ; zhu2020secure, ; le2020adversarial, ; liu2021secure, ; atli2020waffle, ; yang2022watermarking, ). The goal of model watermarking is to verify the ownership of a model by encrypting a specific pattern. For example, white-box watermarking methods assume that the parameters of a suspicious model can be accessed, thus they encode a certain message into the target model’s parameters and decode it with a fixed neural network (uchida2017embedding, ); black-box watermarking methods assume that the model owner can only query the suspicious model without access to its parameters, thus they encode a backdoor into the model by forcing the target model to memorize a special input-output mapping (adi2018turning, ). If a model is suspected to be an illegal copy of the target model, the model owner can check the existence of the watermark to claim the ownership. Model watermarking does minimal harm to the model utility, because it does not interfere with the regular training process. Due to the natural gap between the data structure in computer vision and recommender systems, as well as the gap between the classification task and the ranking task, the watermarking methods in computer vision cannot be applied to recommender systems. Therefore, we aim to fill this gap and propose a model watermarking technique for recommender systems. In this paper, we focus on black-box watermarking, because compared with white-box watermarking, it is more challenging and practical.
@@
There are several unique challenges when designing a black-box watermarking method for recommender systems. First, unlike classification tasks where we can assign a predefined class to the watermark queries, it is non-trivial to define how the output ranking should be like for a watermark query. Since the output is a top-k ranking list, it is possible to encode various forms of watermarks. For example, we can predefine a fixed sub-sequence and insert it in the output, or we can choose a subset of items and force them to always rank higher than another subset, or we can simply promote or demote an item, etc. The best watermark design should take the task-specific constraints and objectives into consideration. Second, the watermark should not exist in an oracle model that is trained only on the regular data. Previous works use out-of-distribution (OOD) data, such as Gaussian noise images (adi2018turning, ) and white pixels (guo2018watermarking, ), to define watermarks. OOD data is not presented in the regular training set, so it is unable to detect watermarks from it. Unfortunately, due to the different data structure (images v.s. sequences), these OOD forms cannot be used for recommender systems. Third, the watermark should be resistant to removal attacks such as distillation and fine-tuning. The attackers aware of the existence of the watermark would probably conduct such attacks to remove it. A valid watermark should remain detectable even after these attacks.
@@
To this end, we propose a model watermarking method for recommender systems, namely Autoregressive Out-of-distribution Watermarking (AOW). We focus on inductive recommender systems since it is more powerful and commonly deployed in real-world scenarios, while they are also greatly threatened by distillation attacks (yue2021black, ). We mainly consider the sequential recommendation scenario. We generate a watermark sequence and train the model to memorize this sequence. The watermark is evaluated by asking the model to predict the next item given the former items. This watermark design is simple and effective to avoid any trivial modifications that may ruin the watermark, such as shuffling and swapping, where the attacker may take such methods to avoid any watermark that potentially exists. We generate the watermark sequence autoregressively. Specifically, we first train an oracle model with the regular training set. Then we select an item as the initial item of the watermark sequence, and query the oracle model with this single item to get the predicted scores for all items. We randomly choose one item from those with lowest scores as the next item of the watermark sequence. We repeat this process until the sequence reaches a predefined length. This autoregressive method can guarantee that the oracle model has zero accuracy on the watermark, as long as the set of items with top-k scores does not overlap with the set of items with lowest scores that we consider for generating the watermark sequence. Besides, AOW is resistant to distillation and fine-tuning as shown in the experiment section. We can still extract the watermark with a high confidence after both attacks. Our code is publicly available [^2] https://github.com/RinneSz/AOW-Autoregressive-Ranking-Optimization.
We summarize our contributions as follows:
@@
  • We propose Autoregressive Out-of-distribution Watermarking (AOW). To the best of our knowledge, this is the first work exploring model watermarking for recommender systems.
@@ 2.1. Model Watermarking
There are plenty of works exploring model watermarking in other domains, such as computer vision (uchida2017embedding, ; chen2019deepmarks, ; wang2020watermarking, ; zhang2018protecting, ; adi2018turning, ; guo2018watermarking, ; zhu2020secure, ; le2020adversarial, ), graph learning (zhao2021watermarking, ; xu2021watermarking, ), federated learning (shao2022fedtracker, ; liang2023fedcip, ; atli2020waffle, ; li2022fedipr, ; liu2021secure, ; yang2022watermarking, ; chen2023fedright, ), etc. They can be divided into white-box and black-box techniques. White-box methods assume that the defender can access the model parameters of a suspicious model, whereas black-box methods assume that the defender can only query the suspicious model and observe the output.
White-box methods usually embed the watermark into model parameters. In computer vision, Uchida et al. (uchida2017embedding, ) jointly train the model with the original objective and the watermarking objective. They reshape the model parameters into a flat vector with size $M$, and multiply it with a key matrix with size $T\times M$, and get a watermarking vector with size $T$, then use a step function on the vector to obtain a vector where each entry is either 1 or 0. With predefined watermarking vector and key matrix, they can learn the model parameters. Successive works (chen2019deepmarks, ; wang2020watermarking, ) mainly follow this design. In federated learning, the goal of model watermarking has shifted to identifying the malicious client, but the basic idea is still to embed the watermark into model parameters (shao2022fedtracker, ; liang2023fedcip, ).
Black-box methods assign a selected label to predefined backdoor patterns such as out-of-distribution data or dedicated noise, and use such data to train the target model, so that when querying the model with these patterns, the model will predict the selected label. Zhang et al. (zhang2018protecting, ) propose to train the target model with OOD/pattern/noise data to embed watermarks. Adi et al. (adi2018turning, ) propose to embed abstract images as watermarks, so it is difficult for the attacker to remove the memorization of those abstract images. Guo et al. (guo2018watermarking, ) propose to generate a pixel pattern according to the secret binary key of the owner, then add the pattern to the clean images to train the model. Zhu et al. (zhu2020secure, ) propose to use a one-way hash function to produce a one-way chain of trigger images. Le Merrer et al. (le2020adversarial, ) propose to generate a set of images that are around the classification boundary of the target model, then fine-tune the model so that those images are well-classified. In graph learning, several works (zhao2021watermarking, ; xu2021watermarking, ) have proposed to use ER random graphs as the watermark and assign predefined labels to the nodes. There are also a large number of works exploring black-box model watermarking on federated learning (atli2020waffle, ; li2022fedipr, ; liu2021secure, ; yang2022watermarking, ; chen2023fedright, ). Their methods basically follow the ideas in computer vision, but innovatively adopted to the federated scenario.
@@ 2.2. Watermarking Against Distillation
Yang et al. (yang2019effectiveness, ) found that the above OOD watermarks can be easily removed by a model distillation. Therefore they propose to use another ingrainer model to help train the target model. The ingrainer model is trained on the OOD watermarks. Then it is fixed, and its output is used to train the target model. An additional loss term is used to let the prediction of the target model approximate the output of the ingrainer model. Li et al. (li2019prove, ) propose to generate real-like samples using a GAN, where both the clean sample and the owner’s watermark logo are the input to an encoder to generate watermarked samples. Then the target model is trained to well classify the watermarked sample and the corresponding benign sample. Namba et al. (namba2019robust, ) propose to generate watermarks by using in-distribution data to counter model stealing attacks, and meanwhile forcing the model to predict the trigger images by depending on model parameters with a large value, so that it will not be pruned by model pruning. The trigger images are sampled from the original training data and are relabeled with a wrong label, such as labeling a dog as a cat. They claim that the model trained under this method will then predict ”dog” for other dog images while predict ”cat” for the trigger image. Jia et al. (jia2021entangled, ) propose to defend against model stealing attacks by adding triggers to in-distribution images. The trigger position is calculated by the largest gradients of the SNNL function in order to avoid bad triggers. Szyller et al. (szyller2021dawn, ) propose to alter the output label of the target model. The selection of altering is conducted by hashing. This method inevitably degrades the utility of the target model.
All these methods cannot be simply transferred to the recommender systems due to the natural gap between the data format as well as the gap between the classification task and the recommendation task.
@@ 3.3. Discussion
White-box watermarks assume that the model owner has access to a suspicious model’s architecture and parameters. In this scenario, the owner can simply encode the watermark in the model parameters, which is already a well-studied technique in previous works (uchida2017embedding, ; chen2019deepmarks, ; wang2020watermarking, ) and can be adopted to recommender systems directly. However, it is not always feasible to access the parameters of a suspicious model. So we focus on black-box watermarking where we can only query the suspicious model and observe its output.
@@
Previous black-box watermark designs in computer vision domain can be categorized into out-of-distribution (OOD) watermarks (zhang2018protecting, ; adi2018turning, ; guo2018watermarking, ) and in-distribution (ID) watermarks (yang2019effectiveness, ; li2019prove, ; namba2019robust, ; szyller2021dawn, ). OOD watermarks are embedded in samples that do not follow the same distribution as the original dataset, while ID watermarks are embedded in samples that follow the original distribution. OOD watermarks do little harm to the model utility, but are not robust against removal attacks. On the contrary, ID watermarks are hard to be removed, but they inevitably degrade the model utility. Since recommender systems are highly correlated to commercial profits, utility degradation is unacceptable in many cases. Therefore, we choose to design OOD watermarks. However, as we will show in the experiment section, AOW is also effective in resisting removal attacks.
@@ 3.4. Autoregressive Out-of-distribution Watermarking (AOW)
To this end, we propose to generate an entire sequence $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{n}\}$ as the watermark. The length $n$ is a hyperparameter to be specified by the model owner. The model should be able to memorize this sequence by predicting the $j$-th item $i^{wm}_{j}$ when given the previous $j-1$ items $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{j-1}\}$. We first train an oracle model using the original dataset, then generate $S_{wm}$ autoregressively. An illustration of AOW is shown in Fig. 1. Specifically, we first select an initial item $i^{wm}_{1}$ (the selection of the initial item will be discussed in the experiment section), then query the oracle model with this single item to get scores of all items. We then rank all items according to the scores and focus on a certain number of items with the lowest scores (let’s say bottom-M in contrast to top-k, where the setting of M will be discussed later on). We draw one item randomly from the bottom-M items, and append it after the initial item to expand the sequence. For model owners who want to design their unique watermark pattern, they can pick one particular item from the bottom-M items by themselves. This may also serve as an effectiveness fingerprinting technique in federated recommendation (yuan2023manipulating, ; yuan2023interaction, ; qu2024towards, ; yuan2023hide, ; zhang2023comprehensive, ) to detect malicious clients who are responsible for the model leakage, which we leave for future work. Now the sequence contains two items, and we query the oracle model again with it and repeat the above process to get the third item. We stop when the sequence reaches a predefined length. This sequence is defined as the watermark sequence $S_{wm}$. It will be used together with the original training set to train a new model $f_{wm}$, and this model will be able to memorize $S_{wm}$ while maintaining its utility. For evaluation, we generate $n-1$ truncated sequences from $S_{wm}$. For example, $S^{1}_{wm}=\{i^{wm}_{1}\}$, $S^{2}_{wm}=\{i^{wm}_{1},i^{wm}_{2}\}$, …, $S^{n-1}_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{n-1}\}$. The model needs to predict the next item given these truncated sequences. A well-memorized watermark should have the ground-truth next item ranked as high as possible.
This design has several advantages.

### v2

@@ Abstract.
Recommender systems embody significant commercial value and represent crucial intellectual property. However, the integrity of these systems is constantly challenged by malicious actors seeking to steal their underlying models. Safeguarding against such threats is paramount to upholding the rights and interests of the model owner. While model watermarking has emerged as a potent defense mechanism in various domains, its direct application to recommender systems remains unexplored and non-trivial. In this paper, we address this gap by introducing Autoregressive Out-of-distribution Watermarking (AOW), a novel technique tailored specifically for recommender systems. Our approach entails selecting an initial item and querying it through the oracle model, followed by the selection of subsequent items with small prediction scores. This iterative process generates a watermark sequence autoregressively, which is then ingrained into the model’s memory through training. To assess the efficacy of the watermark, the model is tasked with predicting the subsequent item given a truncated watermark sequence. Through extensive experimentation and analysis, we demonstrate the superior performance and robust properties of AOW. Notably, our watermarking technique exhibits high-confidence extraction capabilities and maintains effectiveness even in the face of distillation and fine-tuning processes.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679617isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systemsccs: Information systems Data miningccs: Computing methodologies Neural networks
@@ 1. Introduction
Machine learning models have become a necessary component in the modern information society. One of the most representative machine learning products is the recommender system (Rendle et al., 2012; Zhang et al., 2019; He et al., 2020; Wu et al., 2022; Long et al., 2024). It has been widely deployed in various domains including health care (Yue et al., 2021b), finance (Sharaf et al., 2022), e-commerce (Schafer et al., 2001), social platforms (Fan et al., 2019), etc. It plays a crucial role in improving user experience and gaining profits for the service provider. As highly commercialized products, many of them involve certain intellectual property and copyright as well as business secrets. Some recommender systems are developed by the service provider themselves, where cutting-edge techniques might have been adopted to outperform competitors. In this case, model theft and model leakage are intolerable. Some other recommender systems are developed by a third party at the request of the service provider, where the illegal use and redistribution of the model are typically prohibited. Thus, it is necessary to take protective measures to ensure the recommender systems are not being used for these malicious aims.
@@
To the best of our knowledge, there are a limited number of works addressing the above issue on recommender systems. Zhang et al. (Zhang et al., 2024) proposed an optimization-based method against model extraction attacks on recommender systems. They train the target model with an additional loss that can cause malfunctions to the attacker’s model. However, this method inevitably degrades the performance of the target model. As a commercial product, any degradation in performance for the recommender system may lead to a huge profit loss. Thus, it is necessary to design a protection method without sacrificing the model utility.
@@
One promising defense strategy in this scenario is the model watermarking. It has been extensively studied in computer vision (Uchida et al., 2017; Chen et al., 2019; Wang et al., 2020; Zhang et al., 2018; Adi et al., 2018; Guo and Potkonjak, 2018; Zhu et al., 2020; Le Merrer et al., 2020; Liu et al., 2021; Atli et al., 2020; Yang et al., 2022). The goal is to verify the ownership of a model by encrypting a specific pattern. For example, white-box watermarking methods assume that the parameters of a suspicious model can be accessed, thus they encode a certain message into the target model’s parameters and decode it with a fixed neural network (Uchida et al., 2017); black-box watermarking methods assume that the model owner can only query the suspicious model without access to its parameters, thus they encode a backdoor into the model by forcing the target model to memorize a special input-output mapping (Adi et al., 2018). If a model is suspected to be an illegal copy of the target model, the model owner can check the existence of the watermark to claim the ownership. Model watermarking does minimal harm to the model utility, because it does not interfere with the regular training process. Due to the natural gap between the data structure in computer vision and recommender systems, as well as the gap between the classification task and the ranking task, the watermarking methods in computer vision cannot be applied to recommender systems. Therefore, we aim to fill this gap and propose a model watermarking technique for recommender systems. In this paper, we focus on black-box watermarking, because compared with white-box watermarking, it is more challenging and practical.
@@
There are several unique challenges when designing a black-box watermarking method for recommender systems. First, unlike classification tasks where we can assign a predefined class to the watermark queries, it is non-trivial to define how the output ranking should be like for a watermark query. Since the output is a top-k ranking list, it is possible to encode various forms of watermarks. For example, we can predefine a fixed sub-sequence and insert it in the output, or we can choose a subset of items and force them to always rank higher than another subset, or we can simply promote or demote an item, etc. The best watermark design should take the task-specific constraints and objectives into consideration. Second, the watermark should not exist in an oracle model that is trained only on the regular data. Previous works use out-of-distribution (OOD) data, such as Gaussian noise images (Adi et al., 2018) and white pixels (Guo and Potkonjak, 2018), to define watermarks. OOD data is not presented in the regular training set, so it is unable to detect watermarks from it. Unfortunately, due to the different data structure (images v.s. sequences), these OOD forms cannot be used for recommender systems. Third, the watermark should be resistant to removal attacks such as distillation and fine-tuning. The attackers aware of the existence of the watermark would probably conduct such attacks to remove it. A valid watermark should remain detectable even after these attacks.
@@
To this end, we propose a model watermarking method for recommender systems, namely Autoregressive Out-of-distribution Watermarking (AOW). We focus on inductive recommender systems since it is more powerful and commonly deployed in real-world scenarios, while they are also greatly threatened by distillation attacks (Yue et al., 2021a). We mainly consider the sequential recommendation scenario. We generate a watermark sequence and train the model to memorize this sequence. The watermark is evaluated by asking the model to predict the next item given the former items. This watermark design is simple and effective to avoid any trivial modifications that may ruin the watermark, such as shuffling and swapping, where the attacker may take such methods to avoid any watermark that potentially exists. We generate the watermark sequence autoregressively. Specifically, we first train an oracle model with the regular training set. Then we select an item as the initial item of the watermark sequence, and query the oracle model with this single item to get the predicted scores for all items. We randomly choose one item from those with lowest scores as the next item of the watermark sequence. We repeat this process until the sequence reaches a predefined length. This autoregressive method can guarantee that the oracle model has zero accuracy on the watermark, as long as the set of items with top-k scores does not overlap with the set of items with lowest scores that we consider for generating the watermark sequence. Besides, AOW is resistant to distillation and fine-tuning as shown in the experiment section. We can still extract the watermark with a high confidence after both attacks. Our code is publicly available [^1] https://github.com/RinneSz/AOW-Autoregressive-Ranking-Optimization. We summarize our contributions as follows:
@@
  • We propose Autoregressive Out-of-distribution Watermarking (AOW). To the best of our knowledge, it is the first work exploring model watermarking for recommender systems.
@@ 2.1. Model Watermarking
There are plenty of works exploring model watermarking in other domains, such as computer vision (Uchida et al., 2017; Chen et al., 2019; Wang et al., 2020; Zhang et al., 2018; Adi et al., 2018; Guo and Potkonjak, 2018; Zhu et al., 2020; Le Merrer et al., 2020), graph learning (Zhao et al., 2021; Xu et al., 2021), federated learning (Shao et al., 2022; Liang and Wang, 2023; Atli et al., 2020; Li et al., 2022; Liu et al., 2021; Yang et al., 2022; Chen et al., 2023), etc. They can be divided into white-box and black-box techniques. White-box methods assume defenders can access the model parameters of a suspicious model, whereas black-box methods assume defenders can only query the suspicious model and observe the output.
White-box methods usually embed the watermark into model parameters. In computer vision, Uchida et al. (Uchida et al., 2017) encrypts the watermark into the model parameters by multiplying the parameters with a key matrix. Successive works (Chen et al., 2019; Wang et al., 2020) mainly follow this design. In federated learning, the goal of model watermarking has shifted to identifying the malicious client, but the basic idea is still to embed the watermark into model parameters (Shao et al., 2022; Liang and Wang, 2023).
Black-box methods assign a selected label to predefined backdoor patterns such as out-of-distribution data (Zhang et al., 2018; Adi et al., 2018; Zhu et al., 2020) or dedicated noise (Guo and Potkonjak, 2018; Le Merrer et al., 2020), and use such data to train the target model, so that when querying the model with these patterns, the model will predict the selected label. In graph learning, several works (Zhao et al., 2021; Xu et al., 2021) have proposed to use ER random graphs as the watermark and assign predefined labels to the nodes. There are also a large number of works exploring black-box model watermarking on federated learning (Atli et al., 2020; Li et al., 2022; Liu et al., 2021; Yang et al., 2022; Chen et al., 2023). Their methods basically follow the ideas in computer vision, but innovatively adopted to the federated scenario.
@@ 2.2. Watermarking Against Distillation
Yang et al. (Yang et al., 2019) found that the above OOD watermarks can be easily removed by a model distillation. They propose to train an ingrainer model on the watermarks. Then it is fixed as a teacher to train the target model. Successive works mainly focus on generating in-distribution and real-like watermark samples. Li et al. (Li et al., 2019) propose to generate real-like watermark samples using a GAN. Namba et al. (Namba and Sakuma, 2019) propose to generate watermarks by relabeling benign samples as target watermark labels. Jia et al. (Jia et al., 2021) propose to add triggers to in-distribution images. Szyller et al. (Szyller et al., 2021) propose to alter the output label of the target model. The selection of altering is conducted by hashing. This method inevitably degrades the utility of the target model. All these methods cannot be simply transferred to the recommender systems due to the natural gap between the data format as well as the gap between the classification task and the recommendation task.
@@ 3.3. Discussion
White-box watermarks assume that the model owner has access to a suspicious model’s architecture and parameters. In this scenario, the owner can simply encode the watermark in the model parameters, which is already a well-studied technique in previous works (Uchida et al., 2017; Chen et al., 2019; Wang et al., 2020) and can be adopted to recommender systems directly. However, it is not always feasible to access the parameters of a suspicious model. So we focus on black-box watermarking where we can only query the suspicious model and observe its output.
@@
Previous black-box watermark designs in computer vision domain can be categorized into out-of-distribution (OOD) watermarks (Zhang et al., 2018; Adi et al., 2018; Guo and Potkonjak, 2018) and in-distribution (ID) watermarks (Yang et al., 2019; Li et al., 2019; Namba and Sakuma, 2019; Szyller et al., 2021). OOD watermarks are embedded in samples that do not follow the same distribution as the original dataset, while ID watermarks are embedded in samples that follow the original distribution. OOD watermarks do little harm to the model utility, but are not robust against removal attacks. On the contrary, ID watermarks are hard to be removed, but they inevitably degrade the model utility. Since recommender systems are highly correlated to commercial profits, utility degradation is unacceptable in many cases. Therefore, we choose to design OOD watermarks. However, as we will show in the experiment section, AOW is also effective in resisting removal attacks.
@@ 3.4. Autoregressive Out-of-distribution Watermarking (AOW)
To this end, we propose to generate an entire sequence $S_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{n}\}$ as the watermark. The length $n$ is a hyperparameter to be specified by the model owner. The model should be able to memorize this sequence by predicting the $j$-th item $i^{wm}_{j}$ when given the previous $j-1$ items $\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{j-1}\}$. We first train an oracle model using the original dataset, then generate $S_{wm}$ autoregressively. An illustration of AOW is shown in Fig. 1. Specifically, we first select an initial item $i^{wm}_{1}$ (the selection of the initial item will be discussed in the experiment section), then query the oracle model with this single item to get scores of all items. We then rank all items according to the scores and focus on a certain number of items with the lowest scores (let’s say bottom-M in contrast to top-k, where the setting of M will be discussed later on). We draw one item randomly from the bottom-M items, and append it after the initial item to expand the sequence. For model owners who want to design their unique watermark pattern, they can pick one particular item from the bottom-M items by themselves. This may also serve as an effectiveness fingerprinting technique in federated recommendation (Yuan et al., 2023c, a; Qu et al., 2024; Yuan et al., 2023b; Zhang et al., 2023) to detect malicious clients who are responsible for the model leakage, which we leave for future work. Now the sequence contains two items, and we query the oracle model again with it and repeat the above process to get the third item. We stop when the sequence reaches a predefined length. This sequence is defined as the watermark sequence $S_{wm}$. It will be used together with the original training set to train a new model $f_{wm}$, and this model will be able to memorize $S_{wm}$ while maintaining its utility. For evaluation, we generate $n-1$ truncated sequences from $S_{wm}$. For example, $S^{1}_{wm}=\{i^{wm}_{1}\}$, $S^{2}_{wm}=\{i^{wm}_{1},i^{wm}_{2}\}$, …, $S^{n-1}_{wm}=\{i^{wm}_{1},i^{wm}_{2},...,i^{wm}_{n-1}\}$. The model needs to predict the next item given these truncated sequences. A well-memorized watermark should have the ground-truth next item ranked as high as possible. This design has several advantages.